OpenStudy (anonymous):

what are all that justify the following statement. 2 * 1/2= 1

OpenStudy (anonymous):

\[(2/1)*(2 * (1/2)) = 1 * (2/1)\] \[2 = 2\]

OpenStudy (anonymous):

\[2*\frac{1}{2}=1\]\[\frac{2*1}{2}=1\]\[\frac{2}{2}=1\]\[1=1\]